# A Two-Stage GPU Kernel Tuner Combining Semantic Refactoring and Search-Based Optimization

### v1

@@ 4 Experimental Setup
Metrics. Our evaluation considers both correctness and performance. For correctness, we construct test cases with diverse tensor input shapes and compare the outputs of the final kernels optimized by our method against the outputs produced by the original SGLang implementation, which serves as the ground truth [ref39],. For performance, we measure the execution time of the original SGLang kernels, the state-of-the-art Astra variants, and our optimized variants on the same tensor shapes, and report speedup as the primary metric.
Kernels. We evaluate three CUDA kernels extracted from the LLM serving framework SGLang [ref39, ref40]: silu_and_mul, fused_add_rmsnorm, and merge_attn_states_lse. Their computations are summarized in Table LABEL:tab:kernel_comp.
Performance measurement. We benchmark each kernel on representative input shapes and report two types of results: (i) the best configuration for each shape, and (ii) a general configuration that achieves the best average performance across all shapes. For each input shape and each candidate configuration, we perform 20 warm-up runs, then repeat the kernel execution 100 times and report the average kernel execution time. We use SGLang’s measured time as the baseline. The input shapes are drawn from typical scale settings in modern large-model inference, such as common dimensions for LLaMA-7B, 13B, and 70B [ref36].

### v2

@@ 4 Experimental Setup
Evaluation Metrics. Our evaluation considers both correctness and performance. For correctness, we construct a test suite $\mathcal{T}$ covering different data types and input shapes. For each test case, we compare the output of the final kernel produced by our method against the original SGLang implementation element-wise under identical inputs, and deem it correct if the mismatch is within a given tolerance. For performance, we measure the runtime of the native SGLang , the baseline method Astra [wei2025astra], and ours, and use the speedup over SGLang as the primary metric.
Kernels. We select three representative CUDA kernels from the LLM serving framework SGLang [wei2025astra] as evaluation targets, corresponding to silu_and_mul, fused_add_rmsnorm, and merge_attn_states, which cover representative workload characteristics across different inference stages. Experiments are conducted on multiple data types and input shapes:
  1. (1)
    Kernel1: We test 5 shapes $(n,h,d)$ under fp16/bf16: $(512,32,256)$, $(512,40,128)$, $(768,32,256)$, $(768,40,128)$, $(512,64,128)$.
  2. (2)
    Kernel2: We test 8 shapes $(m,n)$ under fp32: $(128,4096)$, $(256,4096)$, $(1024,4096)$, $(2048,8192)$, $(128,11008)$, $(256,13824)$, $(512,14336)$, $(1024,8192)$.
  3. (3)
    Kernel3: We test 5 shapes $(m,n)$ under fp16/bf16: $(16,4096)$, $(32,5120)$, $(32,8192)$, $(16,12288)$, $(64,8192)$.
Performance Measurement. We report two types of results on representative input shapes: the best specialized configuration per shape, and the best general configuration that maximizes the average performance across shapes. For each input shape and candidate configuration, we first perform 20 warm-up runs to eliminate cold-start effects, and then run 100 repetitions and take the average as the kernel runtime. Using SGLang’s measured time as the baseline, we report the relative speedups of Astra and our method. The input shapes are chosen to reflect common dimension configurations in mainstream LLM inference (e.g., typical hidden sizes and head dimensions for LLaMA-7B, 13B, and 70B [wei2025astra]), ensuring representative evaluation.